Suppose a firm faces a cost function of c= 8+4q+q².


a. What is the firm's fixed cost


b. Drive an expression for the firm's average variable fixed cost

1
Expert's answer
2022-02-08T10:48:23-0500

a) The firm's fixed cost:

FC = 8

b) Average variable cost:

="\frac{total variable cost}{Q}"


="\frac{4q+q^{2}}{q}"


"AVC=4+q"


Average fixed cost:

="\frac{totalfixedcost}{q}"


AFC = "\frac{8}{q}"